OpenStudy (anonymous):

Find the slope by solving for y -8x - 5y = 18 Show all steps

OpenStudy (anonymous):

-5y=18+8x y=-18/5-8/5x ur slope is -8/5

OpenStudy (anonymous):

-8x - 5y = 18 +8 x +8 x -5y = 8x + 18 /-5 /-5 y = -8/5x -18/5 SLope: -8/5

OpenStudy (anonymous):

\[ -8x - 5y = 18 \] \[\Rightarrow 5y = -8x -18 \] \[\Rightarrow y = -\frac{8}{5}x -\frac{18}{5} \] Thus slope = \[ -\frac{8}{5} \]

OpenStudy (anonymous):

I got that one right also! :)